OpenStudy (anonymous):

A waiter is carrying a tray above his head and walking at a constant velocity. If he applies a force of 5.0 newtons on the tray and covers a distance of 10.0 meters, how much is the work being done?

OpenStudy (souvik):

\(W=F*s*cos\theta\) where \(W\)=work done \(F\)=force \(s\)=distance and \(theta\)= the angle between \(F\) and \(s\)